OpenStudy (anonymous):

Solve for x. x^2 + 8x + 9 = -7

OpenStudy (anonymous):

Add -7 to both the sides first and tell me what you get...

OpenStudy (lgbasallote):

careful..

OpenStudy (anonymous):

Ha ha ha.. I will..

OpenStudy (lgbasallote):

if you add - 7 to both sides x^2 + 8x + 9 -7 = -7 - 7 -> x^2 + 8x + 2 = -14 you werent able to make it simple ;)

OpenStudy (anonymous):

Ohh.. I will have to be careful.. Simply add 7 to both the sides..